A child's coin bank contains $1.52 in pennies and nickels. If the number of pennies is 32 less than 3
times the number of nickels, how many pennies are in the bank?

Answer by stanbon(75887)

A child's coin bank contains $1.52 in pennies and nickels.
If the number of pennies is 32 less than 3 times the number of nickels,
how many pennies are in the bank?
-------------------------------
Quantity Equation: p = 3n-32
Value Equation:::: p + 5n = 152 cents
-------------------------------------------
Eequation to solve for "n":
3n-32 = 152-5n
8n = 184
n = 23 (# of nickels)
---
Since p = 3n-32, p = 3*23-32 = 69-32 = 37 (# of pennies)
========================================================
Cheers,
Stan H.